OpenStudy (mallorysipp234):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 100 y = x + 2

OpenStudy (anonymous):

\[x ^{2}+\left( x+2 \right)^{2}=100\] \[x ^{2}+ x ^{2}+4x+4-100=0\] \[2x ^{2}+4x-96=0\] \[x ^{2}+2x-48=0\] 1*-48=-48 8-6=2 8*-6=-48 write 2x=8x-6x and make factors.